OpenStudy (anonymous):

Simplify the expression. (-64)^1/3

OpenStudy (anonymous):

it is asking for the cube root of -64

OpenStudy (anonymous):

Not cubed, squared.

OpenStudy (anonymous):

the denominator of the exponent is the root

OpenStudy (anonymous):

I'll just post the actual equation for you so you can see it.

OpenStudy (anonymous):

but you can still tell me what the cube root of -64 is??

OpenStudy (anonymous):

that is the same as \[\sqrt[3]{-64}\]

OpenStudy (anonymous):

what number times itself 3 times is -64 y^3=-64, what is y

OpenStudy (anonymous):

Hero answered your earlier post....... (-64)^1/3 = 3 root (-64) which means find a number that when multiplied by itself three times produces -64. That number is -4, therefore: 3 root(-64) = -4 because (-4)(-4)(-4) = -64

OpenStudy (anonymous):

\[\sqrt[3]{-64}=-4\]